# Check, Locate, Rectify: A Training-Free Layout Calibration System for Text-to-Image Generation

### v1

@@ Abstract
Diffusion models have recently achieved remarkable progress in generating realistic images. However, challenges remain in accurately understanding and synthesizing the layout requirements in the textual prompts. To align the generated image with layout instructions, we present a training-free layout calibration system SimM that intervenes in the generative process on the fly during inference time. Specifically, following a “check-locate-rectify” pipeline, the system first analyses the prompt to generate the target layout and compares it with the intermediate outputs to automatically detect errors. Then, by moving the located activations and making intra- and inter-map adjustments, the rectification process can be performed with negligible computational overhead. To evaluate SimM over a range of layout requirements, we present a benchmark SimMBench that compensates for the lack of superlative spatial relations in existing datasets. And both quantitative and qualitative results demonstrate the effectiveness of the proposed SimM in calibrating the layout inconsistencies.
@@
${}^{\dagger}$qual contribution.*Work done during internship at Alibaba Group.
@@ 1 Introduction
Text-to-image generation [26, 24, 16] has emerged as a promising application of AI-generated content (AIGC), demonstrating the remarkable ability to generate synthetic images from conditional text descriptions. This technology has attracted considerable attention in recent years due to its potential impact on various domains such as image customization [29, 40], 3D content creation [21, 17] and virtual reality [2]. Since achieving high-quality and diverse image generation is challenging, recent advancements have witnessed the rise of diffusion models [13, 27]. Diffusion models employ a sequential generation process that gradually refines the generated images by iteratively conditioning on noise variables. This iterative refinement mechanism allows for an improvement in the fidelity and quality.
@@ 3.2 Ablation Study
Figure 7: Effect of the number of localization steps $T^{loc}$. Initiating layout rectification at an earlier stage enhances the fidelity of images.
@@ 4 Related Work
Layout Control in Diffusion Models. Existing progress fails to fully understand the spatial relations of objects in the free-form textual descriptions and reflect them in the synthesized image, especially for complex scenes. Therefore, jointly conditioning on text and layout has been studied, where layout control signals can be bounding boxes [22, 35], segmentation maps [37, 5], and key points [41]. Several methods extend the Stable Diffusion model by incorporating layout tokens into attention layers [16, 38, 42] or training layout-aware adapters [22]. However, requiring additional training on massive layout-image pairs, these approaches lack flexibility in the base model and may degrade the quality of the generated images. Therefore, recent efforts [4, 20, 35] design loss conditioned on layout constraints to update the noised latent together with denoising. Layout-Guidance [4] computes the loss by applying the energy function on the cross-attention map, Attention-Refocusing [20] constrains both cross-attention and self-attention to “refocus” on the correct regions, and BoxDiff [35] designs inner-box, outer-box, and corner spatial constraints. However, they introduce extra computational cost for gradient update, which affects the speed of generation. In contrast, our system directly modifies the activations to conform to the target for rectification, minimizing the computation overhead.
Layout Generation. Previous layout-to-image studies [42, 4] have largely neglected the discussion on layout generation and heavily relied on users to directly provide accurate layout boxes for objects. However, this necessitates assessing the legality of user input and increases the learning and interaction difficulty for users. Moreover, we have observed a substantial decline in the quality of generated images when the provided boxes are insufficiently accurate. Latest efforts [20, 22] have turned to large language models like GPT-4 [19] by creating appropriate prompting templates to generate layouts, while each API request adds response time and incurs additional costs. In this paper, our system provides a light-weight solution based on dependency parsing following with heuristic rules.
@@ F.1 Latency Comparison for Layout Generation
Since our SimM system presents a new solution for generating the target layout, we provide a brief discussion of the observed increase in latency here. Existing layout-to-image works [20, 22] commonly rely on GPT-4 [19], however, each invocation of the API requires a response time of ~3 seconds. In contrast, thanks to the industrial-strength library, our proposed solution requires an average of only 0.006 seconds for each prompt and does not require a GPU. This significantly improves the user experience for real-time text-to-image generators.
F.2 Comparison with Training-Based Method
@@
F.3 Additional Qualitative Comparison Results
To show the effectiveness of SimM, we illustrate additional qualitative results in Figs. 11 and 12.

### v3

@@ Abstract
Diffusion models have recently achieved remarkable progress in generating realistic images. However, challenges remain in accurately understanding and synthesizing the layout requirements in the textual prompts. To align the generated image with layout instructions, we present a training-free layout calibration system SimM that intervenes in the generative process on the fly during inference time. Specifically, following a “check-locate-rectify” pipeline, the system first analyses the prompt to generate the target layout and compares it with the intermediate outputs to automatically detect errors. Then, by moving the located activations and making intra- and inter-map adjustments, the rectification process can be performed with negligible computational overhead. To evaluate SimM over a range of layout requirements, we present a benchmark SimMBench that compensates for the lack of superlative spatial relations in existing datasets. And both quantitative and qualitative results demonstrate the effectiveness of the proposed SimM in calibrating the layout inconsistencies. Our project page is at https://simm-t2i.github.io/SimM.
@@
[…]
@@ 1 Introduction
Text-to-image generation [31, 29, 20, 11] has emerged as a promising application of AI-generated content (AIGC), demonstrating the remarkable ability to generate synthetic images from conditional text descriptions. This technology has attracted considerable attention in recent years due to its potential impact on various domains such as image customization [34, 45], 3D content creation [26, 22] and virtual reality [4]. Since achieving high-quality and diverse image generation is challenging, recent advancements have witnessed the rise of diffusion models [16, 32]. Diffusion models employ a sequential generation process that gradually refines the generated images by iteratively conditioning on noise variables. This iterative refinement mechanism allows for an improvement in the fidelity and quality.
@@ 3.2 Ablation Study
Figure 7: Effect of the number of localization steps $T^{loc}$. Initiating layout rectification at an earlier stage enhances the fidelity.
@@ 4 Related Work
Layout Control in Diffusion Models. Existing progress fails to fully understand the spatial relations of objects in the free-form textual descriptions and reflect them in the synthesized image, especially for complex scenes. Therefore, jointly conditioning on text and layout has been studied, where layout control signals can be bounding boxes [27, 40], segmentation maps [42, 7, 1], and key points [46]. Several methods extend the Stable Diffusion model by incorporating layout tokens into attention layers [20, 43, 47] or training layout-aware adapters [27]. However, requiring additional training on massive layout-image pairs, these approaches lack flexibility in the base model and may degrade the quality of the generated images. Therefore, recent efforts [6, 25, 40] design loss conditioned on layout constraints to update the noised latent together with denoising. Layout-Guidance [6] computes the loss by applying the energy function on the cross-attention map, Attention-Refocusing [25] constrains both cross-attention and self-attention to “refocus” on the correct regions, and BoxDiff [40] designs inner-box, outer-box, and corner spatial constraints. However, they introduce extra computational cost for gradient update, which affects the speed of generation. In contrast, our system directly modifies the activations to conform to the target for rectification, minimizing the computation overhead.
Layout Generation. Previous layout-to-image studies [47, 6] have largely neglected the discussion on layout generation and heavily relied on users to directly provide accurate layout boxes for objects. However, this necessitates assessing the legality of user input and increases the learning and interaction difficulty for users. Moreover, we have observed a substantial decline in the quality of generated images when the provided boxes are insufficiently accurate. Latest efforts [25, 27, 21] have turned to large language models like GPT-4 [24] by creating appropriate prompting templates to generate layouts, while each API request adds response time and incurs additional costs. In this paper, our system provides a light-weight solution based on dependency parsing following with heuristic rules.
@@ F.1 Latency Comparison for Layout Generation
Since our SimM system presents a new solution for generating the target layout, we provide a brief discussion of the observed increase in latency here. Existing layout-to-image works [25, 27] commonly rely on GPT-4 [24], however, each invocation of the API requires a response time of $\sim$3 seconds. In contrast, thanks to the industrial-strength library, our proposed solution requires an average of only 0.006 seconds for each prompt and does not require a GPU. This significantly improves the user experience for real-time text-to-image generators.
Figure 11: Layout calibration results of images in different styles.
F.2 Generalization Across Diverse Styles
In practical scenarios, users often request the text-to-image generators to produce images in specific styles. In Fig. 11, we show that the stylistic demands for generated images do not hinder the rectification of the layout by SimM.
F.3 Qualitative Results on Other Benchmarks
We additionally present the qualitative results obtained on two latest benchmarks, HRS [2] and TIFA [18], in Fig. 10. These two benchmarks, similar to DrawBench, excessively focus on relative spatial relations. Due to the cost of comprehensive manual evaluation, we take quantitative evaluation on these benchmarks as future work.
F.4 Comparison with Training-Based Method
@@
F.5 Failure Case Analysis
In Fig. 13, we present typical cases of what the human evaluators perceive as errors. The first case is the repeated generation of objects with some in the wrong position. The second case is that multiple objects interact with each other during generation, resulting in incomplete generation. The third case includes missing or unclear objects. These errors are mostly due to the fact that a single adjustment strength parameter $\alpha$ may not be optimal for all generation. This results in insufficient activation enhancement or suppression on the attention map, leading to inaccuracies in generating all objects accurately or preventing the repeated generation.
Figure 13: Typical failure cases identified by human evaluators.
F.6 Additional Qualitative Comparison Results
To show the effectiveness of SimM, we illustrate additional qualitative results in Figs. 14 and 15.